A very small ball has a mass of 5.00 x 1023 kg and a charge of 4.00C. What magnitude electric field directed upward will balance the weight of the ball so that the ball is suspended motionless above the ground?
1
Expert's answer
2020-10-16T11:00:09-0400

The force, "F" due to upward electric field, is given by"Eq"

That is "F=Eq"

At balance point "F=Eq=mg"

"E=\frac{mg}{q}=(\frac{5\times1023\times9.8}{4})" "=12531.75\frac{N}{C}"